OpenStudy (xoxolove456):

The hypotenuse of a right triangle measures 5cm. One leg of the right triangle is 3 longer than the other. What is the length of each leg????

OpenStudy (xoxolove456):

And What is the perimeter of this triangle ?

OpenStudy (alisoliman):

5^2 = 9 x^ + x^2 25= 10 x^2 x^= 2.5 x = +- sqrt(2.5)= +-1.6 because it is about dimesions so neglect the negative value x = 1.6 cm 3x= 4.8 cm

OpenStudy (alisoliman):

5^2 = 9 x^2 + x^2 25= 10 x^2 x^= 2.5 x = +- sqrt(2.5)= +-1.6 because it is about dimesions so neglect the negative value x = 1.6 cm 3x= 4.8 cm